### v1

# RegMixMatch: Optimizing Mixup Utilization in Semi-Supervised Learning

Abstract
Consistency regularization and pseudo-labeling have significantly advanced semi-supervised learning (SSL). Prior works have effectively employed Mixup for consistency regularization in SSL. However, our findings indicate that applying Mixup for consistency regularization may degrade SSL performance by compromising the purity of artificial labels. Moreover, most pseudo-labeling based methods utilize thresholding strategy to exclude low-confidence data, aiming to mitigate confirmation bias; however, this approach limits the utility of unlabeled samples. To address these challenges, we propose RegMixMatch, a novel framework that optimizes the use of Mixup with both high- and low-confidence samples in SSL. First, we introduce semi-supervised RegMixup, which effectively addresses reduced artificial labels purity by using both mixed samples and clean samples for training. Second, we develop a class-aware Mixup technique that integrates information from the top-2 predicted classes into low-confidence samples and their artificial labels, reducing the confirmation bias associated with these samples and enhancing their effective utilization. Experimental results demonstrate that RegMixMatch achieves state-of-the-art performance across various SSL benchmarks.
Introduction
Semi-supervised learning (SSL) aims to leverage a small amount of labeled data to enable the model to extract useful information from a large volume of unlabeled data during training. FixMatch (Sohn et al. 2020) and its variants (Zhang et al. 2021; Xu et al. 2021; Wang et al. 2023; Huang et al. 2023) have demonstrated competitive results in SSL by retaining pseudo-labels (Lee et al. 2013) for high-confidence samples using a threshold strategy combined with consistency regularization. However, these methods heavily rely on data augmentation techniques for consistency regularization, and their pseudo-labeling approach is constrained to high-confidence samples alone. This raises two critical questions: 1) Are existing data augmentation techniques optimal for SSL? 2) Are the low-confidence samples filtered by the threshold truly unusable? At the core of these questions is the challenge of effectively utilizing both high- and low-confidence samples. In this paper, we aim to address these questions by proposing solutions that optimize the use of the Mixup technique (Zhang et al. 2018) in SSL.
(a) Purity and accuracy.
(b) Reliability and top-n accuracy.
Figure 1: Motivating example of RegMixMatch. (a) Mixup’s confidence-reducing behavior undermines SSL. (b) The top-2 accuracy rate across all unlabeled data surpasses the reliability of pseudo labels.
Mixup, which generates augmented samples by interpolating between randomly shuffled data, has been introduced into SSL (Berthelot et al. 2019b, a) to enhance model generalization, yielding significant results. However, recent studies (Wen et al. 2021; Pinto et al. 2022) have revealed that Mixup can induce confidence-reducing behavior because it inherently predicts interpolated labels for every input. We argue that this confidence reduction is problematic in SSL because creating effective artificial labels for unlabeled samples is crucial for training. Methods like sharpening (Berthelot et al. 2019a) and pseudo-labeling (Sohn et al. 2020), which harden softmax prediction probabilities, have been shown to be effective. This suggests that SSL expects to achieve purer artificial labels through low-entropy prediction probabilities, thereby better supporting model training. However, Mixup increases the entropy of prediction probabilities, adding noise to artificial labels and consequently affecting SSL classification performance due to the lack of high-purity labels. As illustrated in Figure 1(a), the purity (the proportion of artificial labels where the highest predicted probability exceeds a threshold, reflecting the model’s prediction confidence) and test accuracy are compared when FixMatch is trained using Empirical Risk Minimization (ERM) (Vapnik 1991) and Mixup methods with pseudo labels. The results indicate that Mixup reduces prediction confidence, decreasing the number of high-purity artificial labels and leading to lower classification performance than ERM, contrasting its success in supervised learning (Zhang et al. 2018). RegMixup (Pinto et al. 2022) addresses this by using Mixup as a regularizer, training with both unmixed and mixed samples to mitigate the high-entropy issue. Inspired by this, and to counter the reduction in artificial label purity caused by Mixup’s high-entropy behavior in SSL, we propose Semi-supervised RegMixup (SRM). SRM combines SSL’s unique pseudo-labeling technique with weak-to-strong consistency regularization, applying RegMixup within SSL. However, pseudo-labeling alone only allows SRM to effectively utilize high-confidence samples. In the following section, we will demonstrate how to effectively utilize low-confidence data.
Pseudo-labeling typically employs a fixed or dynamic threshold (Zhang et al. 2021; Xu et al. 2021; Wang et al. 2023; Wei et al. 2023) to filter out low-confidence samples, retaining pseudo labels only for those with confidence above the threshold. While this strategy fully exploits high-confidence samples, it leaves samples with confidence below the threshold underutilized, as the supervisory information from these samples is generally noisier. However, unlike previous approaches that focus exclusively on leveraging high-confidence samples, we argue that low-confidence samples also hold significant potential. In Figure 1(b), we present the reliability of pseudo labels (i.e., the accuracy of predictions for unlabeled data that surpass the threshold) during FixMatch training, alongside the top-1 and top-2 accuracy rates of the model’s predictions for all unlabeled data. Notably, while the top-1 accuracy is relatively low, the top-2 accuracy exceeds the reliability of pseudo labels. This indicates that by developing methods that incorporate information from the top-2 classes into the samples and their artificial labels, we could enhance the utilization of a broader range of samples, thereby generating more reliable artificial labels even for those with low initial confidence.
To this end, we introduce a heuristic method called Class-Aware Mixup (CAM), designed to effectively utilize low-confidence samples by generating mixed samples for model training. Since artificial labels for low-confidence samples often contain significant noise due to uncertain predictions, mixing these samples with high-confidence samples of the same predicted class helps to reduce noise in the artificial labels, thereby improving their quality for these challenging samples. Our approach, dubbed RegMixMatch, integrates SRM and CAM, both of which leverage Mixup to exploit high- and low-confidence samples, respectively. To validate the effectiveness of this approach, we compare the training time and classification performance of RegMixMatch with previous Mixup-based SSL methods in our experiments, demonstrating significant improvements. The key contributions of our work can be summarized as follows:
  • We discover and verify that Mixup reduces the purity of artificial labels in SSL, ultimately leading to a decline in model performance.
  • We propose SRM, which applies RegMixup to SSL by combining pseudo-labeling and consistency regularization to address the aforementioned issue.
  • We introduce CAM for low-confidence samples, aimed at mitigating confirmation bias and fully exploiting the potential of low-confidence data.
  • We conduct extensive experiments to validate the performance of RegMixMatch, achieving state-of-the-art results in most scenarios.
Related Work
Since the advent of deep learning, SSL has experienced rapid advancement. In this paper, we present an overview of SSL from two key perspectives: consistency regularization and pseudo-labeling. Additionally, we provide a brief outline of data augmentation techniques based on mixing strategies.
Consistency Regularization.
Consistency regularization (Bachman, Alsharif, and Precup 2014) strengthens SSL by enforcing the principle that a classifier should maintain consistent class predictions for an unlabeled example, even after the application of data augmentation. This principle has led to the development of various SSL algorithms based on different augmentation techniques. Traditional SSL methods typically employ softmax or sharpened softmax outputs to supervise perturbed data, thereby achieving consistency regularization. For instance, Temporal Ensembling (Laine and Aila 2017) ensures model consistency by periodically updating the output mean and minimizing the difference in predictions for the same input across different training epochs. Mean Teacher (Tarvainen and Valpola 2017) enhances learning targets by generating a superior teacher model through the exponential moving average of model weights. VAT (Miyato et al. 2018) introduces adversarial perturbations to input data, increasing the model’s robustness. More recently, MixMatch (Berthelot et al. 2019b) and ReMixMatch (Berthelot et al. 2019a) have incorporated the Mixup technique to create augmented samples, thereby boosting generalization performance. ReMixMatch, along with UDA (Xie et al. 2020), also integrates advanced image augmentation techniques such as AutoAugment (Cubuk et al. 2019) and RandAugment (Cubuk et al. 2020), which generate heavily distorted yet semantically intact images, further enhancing classification performance.
Pseudo-Labeling.
Pseudo-labeling (Lee et al. 2013) has been demonstrated to offer superior supervision for augmented samples in SSL. To address the issue of confirmation bias inherent in pseudo-labeling, confidence-based thresholding techniques have been developed to enhance the reliability of pseudo labels. FixMatch (Sohn et al. 2020) employs a fixed threshold to generate pseudo labels from high-confidence predictions on weakly augmented images, ensuring consistency with their strongly augmented counterparts. FlexMatch (Zhang et al. 2021) and FreeMatch (Wang et al. 2023) introduce class-specific thresholds that adapt based on each class’s learning progress. Dash (Xu et al. 2021) dynamically adjusts the threshold during training to refine the filtering of pseudo labels. MPL (Pham et al. 2021) leverages a teacher network to produce adaptive pseudo labels for guiding a student network on unlabeled data. SoftMatch (Chen et al. 2023) tackles the trade-off between pseudo-label quality and quantity through a weighting function. Building on pseudo-labeling, FlatMatch (Huang et al. 2023) minimizes a cross-sharpness measure to ensure consistent learning across labeled and unlabeled datasets, while SequenceMatch (Nguyen 2024) introduces consistency constraints between augmented sample pairs, reducing discrepancies in the model’s predicted distributions across different augmented views. Although pseudo-labeling enhances supervision and overall performance in these methods, its inherent limitations restrict the utilization to primarily high-confidence samples.
In this paper, we employ the SRM, specifically designed for SSL, to achieve consistency regularization by addressing the issue of reduced artificial label purity through the retention of a portion of clean samples for training. Additionally, we explore methods to utilize low-confidence samples through CAM. The differences from previous Mixup-based approaches are detailed in the appendix.
Mixing-Based Data Augmentation.
Mixup (Zhang et al. 2018) generates augmented data by performing a linear interpolation between samples and their corresponding labels, aiming to achieve smoother decision boundaries. Unlike pixel-level Mixup, Manifold Mixup (Verma et al. 2019) applies interpolation within the model’s hidden layers, allowing regularization at different representation levels. CutMix (Yun et al. 2019) combines the principles of Mixup and Cutout (DeVries and Taylor 2017) by cutting a region from one image and pasting it onto another, thereby mixing samples while preserving the local structure of the images. SaliencyMix (Uddin et al. 2021) further improves upon CutMix by detecting salient regions, ensuring that the cropped regions express label-relevant features, thus enhancing performance. ResizeMix (Qin et al. 2020) simplifies the process by resizing the source image to a small patch and pasting it onto another image, effectively retaining essential image information. For a more detailed comparison of different image mixing methods, we recommend reviewing OpenMixup (Li et al. 2023). In this work, ResizeMix is chosen as the primary mixing method, with results based on Mixup presented in the appendix for comparison.
Preliminary: From Mixup To RegMixup
Given a training dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, ERM (Vapnik 1991) uses the empirical data distribution $P_{\delta}(x,y)$ from the training set to approximate the true data distribution $P(x,y)$:
$P_{\delta}(x,y)=\frac{1}{n}\sum_{i=1}^{n}\delta(x=x_{i},y=y_{i}),$ (1)
where $\delta(x=x_{i},y=y_{i})$ denotes the Dirac mass centered at $(x_{i},y_{i})$. The loss function is then minimized over the data distribution $P_{\delta}$. In contrast to ERM, Vicinal Risk Minimization (VRM) (Chapelle et al. 2000) employs strategies to generate new data distributions in the vicinity of each sample, thereby estimating a richer distribution that provides a more informed risk assessment within the neighborhood around each sample. Based on VRM, Mixup (Zhang et al. 2018) achieves a convex combination of samples to approximate the true data distribution, enhancing the model’s generalization performance. The vicinal distribution of Mixup can be represented as
$P_{v}(x,y)=\frac{1}{n}\sum_{i=1}^{n}\delta(x=\bar{x}_{i},y=\bar{y}_{i}),$ (2)
where $\bar{x}_{i}=\lambda x_{i}+(1-\lambda)x_{j}$ and $\bar{y}_{i}=\lambda y_{i}+(1-\lambda)y_{j}$. For each $(x_{i},y_{i})$ feature-target vector, the corresponding $(x_{j},y_{j})$ is drawn at random from the training data. $\lambda$ follows a Beta distribution with parameters $(\alpha,\alpha)$. The larger the $\alpha$, the closer $\lambda$ is to 0.5, resulting in stronger interpolation.
Research on out-of-distribution (OOD) detection by RegMixup (Pinto et al. 2022) reveals that Mixup can induce confidence-reducing behavior, impairing the model’s ability to effectively differentiate between in-distribution and out-of-distribution samples. To counteract this issue, they propose that by using Mixup to generate mixed samples while retaining clean samples for training, the high-entropy behavior can be mitigated, thereby enhancing the model’s robustness to OOD samples. In this context, the approximate data distribution $P_{v}(x,y)$ can be constructed as
$P_{v}(x,y)=\frac{1}{n}\sum_{i=1}^{n}\left[\delta(x=x_{i},y=y_{i})+\delta(x=\bar{x}_{i},y=\bar{y}_{i})\right].$ (3)
Based on this distribution, the following loss function is minimized:
$H(y_{i},p_{\theta}(y\mid x_{i}))+H(\bar{y}_{i},p_{\theta}(y\mid\bar{x}_{i})),$ (4)
where $H(\cdot,\cdot)$ denotes the standard cross-entropy loss, and $p_{\theta}$ represents the softmax output of a neural network parameterized by $\theta$. Later, we will introduce the proposed SRM, which integrates RegMixup into SSL by utilizing pseudo-labeling to address the reduction in artificial label purity caused by Mixup.
RegMixMatch
Figure 2: Overview of RegMixMatch. (a) shows the main idea of SRM. A weakly augmented image is fed into the model to obtain predictions. If the prediction confidence surpasses the threshold $\tau_{c}$, a pseudo label is used to compute consistency loss against the model’s prediction on the corresponding strongly augmented view. If the confidence also exceeds the threshold $\tau_{m}$, its strongly augmented view and pseudo label are mixed with those of another high-confidence sample to implement Mixup. (b) illustrates why CAM works. When a low-confidence sample is mispredicted, CAM converts noise (red part) from artificial label into useful information, while MixUp suffers from it. (c) demonstrates how ResizeMix generates mixed images and labels.
Given a batch of labeled data containing $B$ samples $\mathcal{X}=\{(x_{b}^{l},y_{b}):b\in(1,\ldots,B)\}$ and a batch of unlabeled data containing $\mu B$ samples $\mathcal{U}=\{(x_{b}^{u}):b\in(1,\ldots,\mu B)\}$, $\mu$ is a hyperparameter that determines the relative sizes of $\mathcal{X}$ and $\mathcal{U}$. The supervised loss for the labeled samples in SSL methods is typically constructed as the standard cross-entropy loss between the model’s predictions and the true labels:
$\mathcal{L}_{s}=\frac{1}{B}\sum_{b=1}^{B}H(y_{b},p_{m}(y|x_{b}^{l})),$ (5)
where $p_{m}(y|x)$ is the predicted class distribution produced by the model for input $x$.
Despite the variety of data augmentation methods, they all follow consistency regularization when utilizing unlabeled data: the expectation is that the model outputs similar results for differently perturbed versions of the unlabeled data. Based on this idea, the consistency loss is typically constructed as
$\frac{1}{\mu B}\sum_{b=1}^{\mu B}\ell_{c}(p_{m}(y|\alpha(x_{b}^{u})),p_{m}(y|\alpha(x_{b}^{u}))),$ (6)
where $\ell_{c}$ is the consistency loss, which can be cross-entropy or $\ell$-2 loss. $\alpha(x)$ represents a random augmentation of input $x$, meaning the two terms in Equation 6 can have different values due to the different random augmentations applied.
From RegMixup to Semi-supervised RegMixup
The overall architecture of SRM is depicted in Figure 2(a). In supervised learning, RegMixup employs cross-entropy loss on both clean and mixed samples. However, in SSL, a substantial portion of the data remains unlabeled. By leveraging pseudo-labeling, we retain pseudo labels for unlabeled samples that exceed a specific confidence threshold. In this section, we will illustrate how the two terms in Equation 4 are adapted for SSL.
For an unlabeled image $x_{b}^{u}$, we first apply a weak augmentation $\alpha(x_{b}^{u})$ and compute the model’s predicted class distribution $q_{b}=p_{m}(y|\alpha(x_{b}^{u}))$. The class with the highest probability, denoted as $\hat{q}_{b}=\arg\max(q_{b})$, is then assigned as the pseudo label for the corresponding strongly augmented view $\mathcal{A}(x_{b}^{u})$. Consequently, we transform the RegMixup loss function for clean samples into the following form:
$\mathcal{L}_{u}=\frac{1}{\mu B}\sum_{b=1}^{\mu B}\mathds{1}(\max{(q_{b})}\geq\tau_{c})H(\hat{q}_{b},p_{m}(y|\mathcal{A}(x_{b}^{u}))),$ (7)
where $\tau_{c}$ is the threshold above which a pseudo-label is retained, and $\max{(q_{b})}$ represent the maximum predicted probability of the weak augmentation $\alpha(x_{b}^{u})$.
To implement Mixup, we also apply thresholding to retain pseudo labels only for high-confidence unlabeled images. Specifically, for a weakly augmented view $\alpha(x_{i}^{u})$ of an unlabeled image within a data batch, we retain the pseudo label $\hat{q}_{i}$ only if its confidence exceeds the threshold $\tau_{m}$. High-confidence images from this batch are then grouped into a set denoted as =$\mathcal{H}=\{x_{b}^{u}|\max(q_{b})>\tau_{m}\}$. For the loss of mixed data, as shown in Figure 2(a), given a strongly augmented view $\mathcal{A}(x_{i}^{u})$ and its pseudo label $\hat{q}_{i}$ from the set $\mathcal{H}$, we randomly select $\mathcal{A}(x_{j}^{u})$ and $\hat{q}_{j}$ from the same set $\mathcal{H}$, generating mixed images and labels using ResizeMix technique for the calculation of $H(\cdot,\cdot)$. Thus, we transform the loss function of RegMixup for mixed samples into
$\mathcal{L}_{m}=\frac{1}{|\mathcal{H}|}\sum_{i=1}^{|\mathcal{H}|}H(\hat{q}_{i}\oplus\hat{q}_{j},p_{m}(y|\mathcal{A}(x_{i}^{u})\oplus\mathcal{A}(x_{j}^{u}))),$ (8)
where the symbol $\oplus$ denotes the image or label mixing operation, and ||$\cdot$|| ⋅ | represents the number of elements in the set.
Note that SRM employs two different thresholds, $\tau_{c}$ and $\tau_{m}$. It has been observed that when $\tau_{m}$ is close to or lower than $\tau_{c}$, the model is prone to significant confirmation bias, particularly during the early training stages. To mitigate this, $\tau_{m}$ is set to a higher value. Additionally, as described in the preliminary, $\alpha$ is used to control the mixing intensity; a higher $\alpha$ results in a stronger mixing intensity. Our parameter $\alpha_{h}$, which controls the mixing intensity of high-confidence samples, can be set to a higher value than previous Mixup-based SSL methods due to the retention of clean samples, allowing for better generalization performance.
In summary, we select high-confidence samples as clean samples using a high threshold $\tau_{c}$ and retain their pseudo labels to compute the consistency loss. Concurrently, we set an even higher threshold $\tau_{m}$ to filter and retain the high-confidence samples and their pseudo labels for implementing Mixup. Consequently, we can transform Equation 4 into the following form to achieve the transition from RegMixup to SRM:
$\mathcal{L}_{rm}=\mathcal{L}_{u}+\mathcal{L}_{m}.$ (9)
Class-Aware Mixup
Pseudo-labeling enables our SRM to effectively leverage high-confidence samples. However, low-confidence samples are often discarded due to the uncertainty in their predictions. Our proposed CAM method facilitates the utilization of these low-confidence samples while maintaining the quality of the artificial labels.
Figure 2(b) provides a conceptual overview of our CAM. Mixup randomly selects a target image, which can belong to any class, to blend with the source image. For low-confidence unlabeled samples, where the predicted class often diverges from the true class, randomly selecting a target image can lead to confirmation bias. This is because the mixed label may incorporate noisy features from the predicted class, while the mixed image does not reflect these features. However, by limiting the samples mixed with low-confidence ones to high-confidence samples sharing the same predicted classes (i.e., class-aware), the noise can be reduced due to the inclusion of predicted class features. Additionally, Figure 1(b) shows that the top-2 accuracy is high, indicating a strong likelihood that the correct class is among the top-2 predicted classes in the mixed labels, making the inclusion of predicted class information sufficient.
We denote the complement of $\mathcal{H}$ as $\mathcal{H}^{c}$. Given a strongly augmented view $\mathcal{A}(x_{i}^{u})$ of a low-confidence data in $\mathcal{H}^{c}$ and its softmax output $q_{i}$, we use the strongly augmented view $\mathcal{A}(x_{j}^{u})$ and its pseudo label $\hat{q}_{j}$ of the class-aware sample ($\hat{q}_{j}=\arg\max(q_{i})$) from the high-confidence set $\mathcal{H}$ , generating mixed images and labels using ResizeMix technique for the calculation of $\ell$-2 loss (as shown in Equation 10).
$\mathcal{L}_{cm}=\frac{1}{|\mathcal{H}^{c}|}\sum_{i=1}^{|\mathcal{H}^{c}|}\left\|q_{i}\oplus\hat{q}_{j}-p_{m}(y|\mathcal{A}(x_{i}^{u})\oplus\mathcal{A}(x_{j}^{u}))\right\|_{2}^{2}.$ (10)
To reduce the negative impact of uncertain predictions, we use softmax outputs as artificial labels for low-confidence data and apply $\ell$-2 loss instead of cross-entropy loss. Importantly, the parameter $\alpha_{l}$ in CAM, which controls the mixing intensity for low-confidence samples, is set higher than $\alpha_{h}$. This choice helps to minimize the risk of the model overfitting to specific noisy samples, a point that will be further explored in the sensitivity analysis. Although the top-$n$ ($n\geq 3$) accuracy of the model’s predictions would be higher, our experiments show that mixing more than two samples does not yield additional improvements. Ultimately, CAM enables the incorporation of information from the top-2 classes into the samples and their artificial labels, thereby facilitating the utilization of low-confidence samples.
Algorithm 1 RegMixMatch Algorithm
1:  Input: Labeled batch $\mathcal{X}=\{(x_{b}^{l},y_{b}):b\in(1,\dots,B)\}$, unlabeled batch $\mathcal{U}=\{x_{b}^{u}:b\in(1,\dots,\mu B)\}$, confidence threshold $\tau_{c}$ and $\tau_{m}$, mixing intensities $\alpha_{h}$ and $\alpha_{l}$ for high- and low-confidence data
2:  Calculate $\mathcal{L}_{s}$ using Equation 5 $\triangleright$ Loss for labeled data
3:  for $b=1$ to $\mu B$ do
4:     $q_{b}=p_{m}(y|\alpha(x_{b}^{u});\theta)$triangleright Prediction after weak augmentation
5:     $\hat{q}_{b}=\arg\max(q_{b})$triangleright Pseudo labels for unlabeled data
6:  end for
7:  Calculate $\mathcal{L}_{u}$ using Equation 7 $\triangleright$ Loss for clean unlabeled data
8:  Construct high-confidence set =$\mathcal{H}=\{x_{b}^{u}|\max(q_{b})>\tau_{m}\}$
9:  Construct low-confidence set $\mathcal{H}^{c}=\mathcal{U}\setminus\mathcal{H}$
10:  Calculate $\mathcal{L}_{m}$ using Equation 8 $\triangleright$ Loss for mixed high-confidence unlabeled data
11:  $\mathcal{L}_{rm}=\mathcal{L}_{m}+\mathcal{L}_{u}$triangleright Using Mixup as a regularizer
12:  Calculate $\mathcal{L}_{cm}$ using Equation 10 $\triangleright$ Loss for mixed low-confidence unlabeled data
13:  return  $\mathcal{L}_{s}+\mathcal{L}_{rm}+\mathcal{L}_{cm}$
Overall, our proposed SRM employs pseudo-labeling and consistency regularization for high-confidence data, applying Mixup to generate mixed samples while preserving unmixed samples for training in SSL. SRM enhances model generalization by leveraging Mixup, while simultaneously addressing the issue of degraded artificial label purity due to Mixup’s high-entropy behavior. For low-confidence samples, our CAM method mixes them with samples from specific classes, effectively utilizing these challenging samples while mitigating confirmation bias. The full algorithm of RegMixMatch is shown in Algorithm 1. The overall loss of RegMixMatch is formulated as
$\mathcal{L}=\mathcal{L}_{s}+\mathcal{L}_{rm}+\mathcal{L}_{cm}.$ (11)
Dataset CIFAR10 CIFAR100 SVHN STL10
# Label 10 40 250 4000 400 2500 10000 40 250 1000 40 1000
VAT (Miyato et al. 2018) 79.81$\pm$1.17 74.66$\pm$2.12 41.03$\pm$1.79 10.51$\pm$0.12 85.20$\pm$1.40 46.84$\pm$0.79 32.14$\pm$0.19 74.75$\pm$3.38 4.33$\pm$0.12 4.11$\pm$0.20 74.74$\pm$0.38 37.95$\pm$1.12
Mean Teacher (Tarvainen and Valpola 2017) 76.37$\pm$0.44 70.09$\pm$1.60 37.46$\pm$3.30 8.10$\pm$0.21 81.11$\pm$1.44 45.17$\pm$1.06 31.75$\pm$0.23 36.09$\pm$3.98 3.45$\pm$0.03 3.27$\pm$0.05 71.72$\pm$1.45 33.90$\pm$1.37
MixMatch (Berthelot et al. 2019b) 65.76$\pm$7.06 36.19$\pm$6.48 13.63$\pm$0.59 6.66$\pm$0.26 67.59$\pm$0.66 39.76$\pm$0.48 27.78$\pm$0.29 30.60$\pm$8.39 4.56$\pm$0.32 3.69$\pm$0.37 54.93$\pm$0.96 21.70$\pm$0.68
ReMixMatch (Berthelot et al. 2019a) 20.77$\pm$7.48 9.88$\pm$1.03 6.30$\pm$0.05 4.84$\pm$0.01 42.75$\pm$1.05 26.03$\pm$0.35 20.02$\pm$0.27 24.04$\pm$9.13 6.36$\pm$0.22 5.16$\pm$0.31 32.12$\pm$6.24 6.74$\pm$0.14
UDA (Xie et al. 2020) 34.53$\pm$10.69 10.62$\pm$3.75 5.16$\pm$0.06 4.29$\pm$0.07 46.39$\pm$1.59 27.73$\pm$0.21 22.49$\pm$0.23 5.12$\pm$4.27 1.92$\pm$0.05 1.89$\pm$0.01 37.42$\pm$8.44 6.64$\pm$0.17
FixMatch (Sohn et al. 2020) 24.79$\pm$7.65 7.47$\pm$0.28 4.86$\pm$0.05 4.21$\pm$0.08 46.42$\pm$0.82 28.03$\pm$0.16 22.20$\pm$0.12 3.81$\pm$1.18 2.02$\pm$0.02 1.96$\pm$0.03 35.97$\pm$4.14 6.25$\pm$0.33
Dash (Xu et al. 2021) 27.28$\pm$14.09 8.93$\pm$3.11 5.16$\pm$0.23 4.36$\pm$0.11 44.82$\pm$0.96 27.15$\pm$0.22 21.88$\pm$0.07 2.19$\pm$0.18 2.04$\pm$0.02 1.97$\pm$0.01 34.52$\pm$4.30 6.39$\pm$0.56
MPL (Pham et al. 2021) 23.55$\pm$6.01 6.62$\pm$0.91 5.76$\pm$0.24 4.55$\pm$0.04 46.26$\pm$1.84 27.71$\pm$0.19 21.74$\pm$0.09 9.33$\pm$8.02 2.29$\pm$0.04 2.28$\pm$0.02 35.76$\pm$4.83 6.66$\pm$0.00
FlexMatch (Zhang et al. 2021) 13.85$\pm$12.04 4.97$\pm$0.06 4.98$\pm$0.09 4.19$\pm$0.01 39.94$\pm$1.62 26.49$\pm$0.20 21.90$\pm$0.15 8.19$\pm$3.20 6.59$\pm$2.29 6.72$\pm$0.30 29.15$\pm$4.16 5.77$\pm$0.18
SoftMatch (Chen et al. 2023) - 4.91$\pm$0.12 4.82$\pm$0.09 4.04$\pm$0.02 37.10$\pm$0.77 26.66$\pm$0.25 22.03$\pm$0.03 2.33$\pm$0.25 - 2.01$\pm$0.01 21.42$\pm$3.48 5.73$\pm$0.24
FreeMatch (Wang et al. 2023) 8.07$\pm$4.24 4.90$\pm$0.04 4.88$\pm$0.18 4.10$\pm$0.02 37.98$\pm$0.42 26.47$\pm$0.20 21.68$\pm$0.03 1.97$\pm$0.02 1.97$\pm$0.01 1.96$\pm$0.03 15.56$\pm$0.55 5.63$\pm$0.15
SequenceMatch (Nguyen 2024) - 4.80$\pm$0.01 4.75$\pm$0.05 4.15$\pm$0.01 37.86$\pm$1.07 25.99$\pm$0.22 20.10$\pm$0.04 1.96$\pm$0.23 1.89$\pm$0.31 1.79$\pm$0.02 15.45$\pm$1.40 5.56$\pm$0.35
FlatMatch (Huang et al. 2023) 15.23$\pm$8.67 5.58$\pm$2.36 4.22$\pm$1.14 3.61$\pm$0.49 38.76$\pm$1.62 25.38$\pm$0.85 19.01$\pm$0.43 2.46$\pm$0.06 1.43$\pm$0.05 1.41$\pm$0.04 16.20$\pm$4.34 4.82$\pm$1.21
RegMixMatch 4.35$\pm$0.08 4.24$\pm$0.02 4.21$\pm$0.02 3.38$\pm$0.05 35.27$\pm$0.60 23.78$\pm$0.29 19.41$\pm$0.10 1.81$\pm$0.03 1.77$\pm$0.01 1.79$\pm$0.02 11.74$\pm$0.57 4.66$\pm$0.12
Fully-Supervised 4.62$\pm$0.05 19.30$\pm$0.09 2.13$\pm$0.01 -
Table 1: Error rates on CIFAR10/100, SVHN, and STL10 datasets. The fully-supervised results of STL10 are unavailable since we do not have label information for its unlabeled data. The best results are highlighted with Bold and the second-best results are highlighted with underline.
Experiments
In this section, we present an extensive experimental evaluation of the proposed RegMixMatch method. We assess its performance across a variety of widely-used SSL datasets, including CIFAR-10/100 (Krizhevsky, Hinton et al. 2009), SVHN (Netzer et al. 2011), STL-10 (Coates, Ng, and Lee 2011), and ImageNet (Deng et al. 2009), under different labeled data conditions. The experimental results are benchmarked against 13 established SSL algorithms, including VAT (Miyato et al. 2018), Mean Teacher (Tarvainen and Valpola 2017), MixMatch (Berthelot et al. 2019b), ReMixMatch (Berthelot et al. 2019a), UDA (Xie et al. 2020), FixMatch (Sohn et al. 2020), Dash (Xu et al. 2021), MPL (Pham et al. 2021), FlexMatch (Zhang et al. 2021), SoftMatch (Chen et al. 2023), FreeMatch (Wang et al. 2023), SequenceMatch (Nguyen 2024), and FlatMatch (Huang et al. 2023). To underscore the rationale behind leveraging Mixup and the effectiveness of RegMixMatch, we compare its training time and classification performance with prior Mixup-based SSL methods (Berthelot et al. 2019a) and the latest state-of-the-art SSL approaches (Wang et al. 2023; Huang et al. 2023). Additionally, we perform comprehensive ablation studies and hyperparameter analysis to validate the design choices behind RegMixMatch. We also present the results of RegMixMatch on pre-trained backbones in the appendix.
For a fair comparison, and in line with previous SSL studies, we use Wide ResNet-28-2 (Zagoruyko and Komodakis 2016) for CIFAR-10 and SVHN, Wide ResNet-28-8 for CIFAR-100, and ResNet-37-2 (He et al. 2016) for STL-10. All hyperparameter settings are kept consistent with those used in prior work, as detailed in the appendix. For the implementation of RegMixMatch, the threshold $\tau_{c}$ for the consistency loss is set in accordance with the FreeMatch approach, as adopted in the state-of-the-art FlatMatch method. Additionally, we provide experimental results based on the settings of FixMatch in the appendix. The values of $\tau_{m}$, $\alpha_{h}$, and $\alpha_{l}$ are set to 0.999, 1.0, and 16.0, respectively.
Main Results
Table 1 compares the performance of RegMixMatch with existing methods across various datasets. In 12 commonly used SSL scenarios, RegMixMatch achieves state-of-the-art results in 9 cases, demonstrating significant improvement. In the remaining 3 cases, it secures the second-best performance. Notably, previous methods either exhibit minimal improvement or perform poorly in specific scenarios, such as certain datasets or particular labeled data quantities. Our experimental results highlight the comprehensiveness of RegMixMatch. Specifically, RegMixMatch achieves an error rate of only 4.35% on CIFAR-10 with 10 labels and 11.74% on STL-10 with 40 labels, surpassing the second-best results by 3.72% and 3.71%, respectively.
# Label FixMatch FlexMatch FreeMatch RegMixMatch
1w 46.39 45.79 45.31 41.65
10w 28.47 27.83 27.43 26.34
Table 2: Error rates on ImageNet.
Table 2 presents the results of training RegMixMatch on the ImageNet dataset using MAE pre-trained ViT-B (He et al. 2022). RegMixMatch notably outperforms other methods, particularly achieving a 3.66% improvement on ImageNet with 10,000 labels. More comprehensive results and hyperparameter settings are detailed in the appendix.
Efficiency Study
Figure 3: Efficiency analysis of RegMixMatch.
For algorithms handling classification tasks, both runtime and classification performance are critical factors. In this section, we compare RegMixMatch with several SSL algorithms in terms of both metrics. Specifically, we select FlatMatch (the previous state-of-the-art method), FreeMatch (benchmark), and ReMixMatch (Mixup-based SSL method) for comparison. The experiments are conducted on 2 24GB RTX 3090 GPU. Figure 3 presents a comparison of the 4 algorithms under two settings: CIFAR-100 with 2500 labels and STL-10 with 40 labels. The results show that RegMixMatch not only significantly improves classification performance compared to FlatMatch but also requires less training time, highlighting the efficiency of RegMixMatch. Moreover, compared to ReMixMatch and FreeMatch, RegMixMatch achieves substantial performance gains with only a minimal increase in runtime, demonstrating the effectiveness of leveraging Mixup.
Ablation Study
In addition to utilizing low-confidence samples, a key distinction of RegMixMatch from other Mixup-based SSL methods is its retention of clean samples within SRM during training. To verify the rationale behind this design, we compare RegMixMatch with two scenarios: one where only clean samples are retained during training (w/o mixed samples) and another where only mixed samples are used (w/o clean samples). To further validate the rationale and effectiveness of CAM in leveraging low-confidence data, we also present results for three additional configurations: removing CAM entirely (w/o CAM), replacing CAM with Mixup (CAM $\rightarrow$ Mixup) and adapting the CAM to incorporate information from the top-3 predicted classes (Top-2 $\rightarrow$ Top-3). The results for these ablation studies on STL-10 with 40 labels are shown in Table 3.
(a) Purity of artificial labels.
(b) Learning efficiency.
Figure 4: Ablation study of RegMixMatch. (a) demonstrates that RegMixMatch alleviates the reduced purity caused by Mixup. (b) depicts CAM improves learning efficiency.
Table 3 highlights several key findings: 1) Training with only clean samples or only mixed samples yields suboptimal results compared to using both. RegMixMatch leverages SRM not only to enhance generalization performance but also to mitigate the negative effects of high-entropy behavior in SSL, as shown in Figure 4(a). 2) CAM improves both classification accuracy and learning efficiency, as evidenced by the higher accuracy at the same iterations in Figure 4(b). This suggests that CAM facilitates the safer utilization of low-confidence samples, enriching the training process with more informative data. 3) Replacing CAM with Mixup or incorporating features from more specific classes (i.e., Top-3) results in performance degradation. This underscores the importance of specifically limiting the classes and quantities of mixed samples for low-confidence data and demonstrates the rationale of CAM.
Ablation Avg. Error Rate (%)
RegMixMatch 11.74
w/o mixed samples 15.56 (+3.82)
w/o clean samples 15.78 (+4.04)
w/o CAM 12.30 (+0.56)
CAM $\rightarrow$ Mixup 12.12 (+0.38)
Top-2 $\rightarrow$ Top-3 11.96 (+0.22)
Table 3: Ablation results of RegMixMatch.
Sensitivity Analysis
RegMixMatch incorporates three key hyperparameters: the threshold $\tau_{m}$, which retains high-confidence samples for SRM; the parameter $\alpha_{h}$, which controls the mixing intensity for high-confidence samples; and the parameter $\alpha_{l}$, which controls the mixing intensity for low-confidence samples. To assess the specific impact of these hyperparameters on model performance, we conduct experiments on CIFAR-10 with 250 labels.
Confidence Threshold.
As illustrated in Figure 5(a), the optimal value for $\tau_{m}$ is 0.999, which is higher than the threshold $\tau_{c}$ (0.95). During model training, high-confidence clean samples are used to compute the consistency loss based on pseudo labels, validating the model’s predictions on these samples. If $\tau_{m}$ is set to a lower value (e.g., 0.95), it would further validate the model’s predictions on these samples, leading to overconfidence and exacerbating confirmation bias. Thus, $\tau_{m}=0.999$ is chosen to balance the quality and quantity of artificial labels for mixed samples.
Mixing Intensity.
In addition to retaining clean samples, RegMixMatch distinguishes itself from previous Mixup-based SSL methods by generating more robust mixed samples to enhance model generalization. The parameter $\alpha$ controls the mixing intensity, typically set to a value less than 1.0 in prior work. However, as shown in Figures 5(b) and 5(c), the optimal values for $\alpha_{h}$ and $\alpha_{l}$ in RegMixMatch are 1.0 and 16.0, respectively. A higher $\alpha$ indicates stronger augmentation, improving generalization performance. It is noteworthy that $\alpha_{l}\gg\alpha_{h}$. As observed with Mixup, a greater mixing intensity for noisy data enhances the model’s robustness to noise. Since the supervisory information for low-confidence samples is more prone to noise, a higher $\alpha_{l}$ generates more intermediate-state samples, increasing the quantity and variability of the training data, and thereby reducing the risk of overfitting to specific noisy samples.
(a) Threshold $\tau_{m}$.
(b) Mixing intensity $\alpha_{h}$.
(c) Mixing intensity $\alpha_{l}$.
Figure 5: Parameter sensitivity analysis of RegMixMatch.
Conclusion
In this paper, we first demonstrate that Mixup’s high-entropy behavior degrades the purity of artificial labels in SSL. To address this issue, we introduce SRM, a framework that integrates pseudo-labeling with consistency regularization through the application of RegMixup in SSL. We show that SRM effectively mitigates the reduced artificial labels purity, enhancing Mixup’s utility for high-confidence samples. Additionally, we investigate strategies for leveraging low-confidence samples. Specifically, we propose CAM, a method that combines low-confidence samples with specific class samples to enhance the quality and utility of artificial labels. Extensive experiments validate the efficiency and effectiveness of RegMixMatch. Our approach represents an initial exploration into the potential of low-confidence samples, with further in-depth studies recommended as a direction for future research.
Acknowledgments
This work was supported by the Fundamental Research Funds for the Central Universities (No.2022JBMC011).
Appendix for
“RegMixMatch: Optimizing Mixup Utilization in Semi-Supervised Learning”
In this appendix, we provide experimental details of RegMixMatch for reproduction. Additionally, we include supplementary experimental results to verify the superiority of RegMixMatch in terms of artificial label purity and learning efficiency. Finally, we demonstrate the performance of RegMixMatch in two scenarios: large-scale dataset, and pre-trained backbones.
Appendix A Experimental Details
In this section, we present detailed experimental settings and our modifications to ResizeMix (Qin et al. 2020). We also demonstrate the key differences between RegMixMatch and other Mixup (Zhang et al. 2018) based semi-supervised learning methods (Berthelot et al. 2019b, a).
Experimental Setting
Followed FreeMatch (Wang et al. 2023), we categorize the hyperparameter settings into algorithm-dependent and algorithm-independent groups, and present them in Table 4 and Table 5, respectively.
Algorithm RegMixMatch
Unlabeled data to labeled data ratio 7
Threshold $\tau_{m}$ for implementing Mixup 0.999
Mixing intensity $\alpha_{h}$ for high-confidence data 1.0
Mixing intensity $\alpha_{l}$ for low-confidence data 16.0
Thresholding EMA decay 0.999
Table 4: Algorithm-dependent hyperparameters.
Dataset CIFAR-10 CIFAR-100 STL-10 SVHN
Model WRN-28 WRN-28 WRN-37 WRN-28
-2 -8 -2 -2
Weight decay 5e-4 1e-3 5e-4 5e-4
Batch size 64
Learning rate 0.03
SGD momentum 0.9
EMA decay 0.999
Table 5: Algorithm-independent hyperparameters.
ResizeMix
For reproduction, we show our modifications to ResizeMix. ResizeMix resizes the source image through a scaling factor $\tau$ to a smaller patch $P$, where $\tau\sim U(0.1,0.8)$. The resized patch is then pasted onto a random region of the target image. In this way, the mixing ratio $\lambda$ can be expressed using the area ratio of the patch to the target image as follows:
$\lambda=\frac{W_{p}\times H_{p}}{W\times H},$ (12)
where $W$, $H$, $W_{p}$, and $H_{p}$ represent the width and height of the target image and the source patch, respectively. We can observe that $\lambda$ and $\tau$ satisfy
$\lambda=\tau^{2}.$ (13)
To adjust the average mixing intensity, we change the distribution of $\tau$ from uniform to a Beta distribution with parameters $(\alpha,\alpha)$. Then, we use the square root of $\tau$ as the scaling factor to resize the images. Consequently, the mixing ratio $\lambda$ follows a Beta distribution with parameters $(\alpha,\alpha)$. The larger the $\alpha$, the closer $\lambda$ is to 0.5, resulting in stronger mixing. Finally, similar to Mixup, we can control the mixing intensity in ResizeMix using the $\alpha$ parameter.
Key Improvements
MixMatch (Berthelot et al. 2019b) and ReMixMatch (Berthelot et al. 2019a) both use Mixup technique to achieve consistency regularization. These methods generate artificial labels using the softmax (or sharpened softmax) outputs for all unlabeled samples, and they are trained using $\ell$-2 loss. To better distinguish these methods from RegMixMatch, we list the key improvements RegMixMatch introduces when employing Mixup:
  • We categorize data based on confidence levels, with distinct processing methods applied to each category.
  • For high-confidence data, training is done using pseudo labels and cross-entropy loss, while for low-confidence data, the softmax outputs and $\ell$-2 loss are used.
  • The original, non-Mixup samples are retained for training to alleviate the issue of artificial labels purity degradation caused by Mixup.
  • RegMixMatch allows for stronger mixing intensities: the mixing intensity is set to 1.0 for high-confidence data and 16.0 for low-confidence data. The mixing intensity in the other two methods is typically less than 1.0.
Appendix B Additional Experiment Results
In this section, we compare the performance of RegMixMatch with other SSL algorithms in terms of purity and learning efficiency. Additionally, we present the results of applying other thresholding and mixing strategy to RegMixMatch.
Purity Comparison
In Figure 6 we presented the changes in purity during the training process of different SSL algorithms. We can observe that although ReMixMatch enhances the purity of artificial labels by sharpening the softmax outputs, the introduction of Mixup technique results in consistently lower and more fluctuating purity of the artificial labels during training. FreeMatch (Wang et al. 2023) significantly improves the purity of artificial labels by utilizing clean samples for training through the pseudo-labeling technique. Our method, RegMixMatch, while incorporating the Mixup technique, retains the training on clean samples, effectively mitigating the issue of reduced label purity caused by Mixup and achieving nearly the same level of purity as FreeMatch.
Figure 6: Purity of different algorithms.
Figure 7: Learning efficiency of different algorithms.
Learning Efficiency Comparison
We would like to emphasize the superior learning efficiency of RegMixMatch. We presented the changes in accuracy of the RegMixMatch, FixMatch (Sohn et al. 2020), and FreeMatch algorithms during the training process on CIFAR-10 with 250 labels. From Figure 7, we can observe that by effectively leveraging both high and low-confidence samples, RegMixMatch achieves higher accuracy compared to other SSL algorithms within the same number of iterations.
Quantitative Results
As introduced in the paper, RegMixMatch adopts dynamic threshold strategy proposed by FreeMatch for the setting of $\tau_{c}$ and uses ResizeMix technique to generate mixed images. We conduct additional experiments on 1) using fixed threshold strategy (Sohn et al. 2020) for $\tau_{c}$. 2) using Mixup technique for image mixing. The results for these studies on CIFAR-10 are shown in Table 6 and Table 7. We can see that our method demonstrates superior classification performance under both threshold strategies, and the combination of the dynamic threshold strategy with the ResizeMix image augmentation technique achieves the optimal results.
Dataset CIFAR10
# Label 40 250 4000
FixMatch 7.47$\pm$0.28 4.86$\pm$0.05 4.21$\pm$0.08
RegMixMatch (Fix) 6.02$\pm$0.10 4.21$\pm$0.03 3.73$\pm$0.05
FreeMatch 4.90$\pm$0.04 4.88$\pm$0.18 4.10$\pm$0.02
RegMixMatch (Free) 4.24$\pm$0.02 4.21$\pm$0.02 3.38$\pm$0.05
Table 6: Performance on other thresholding strategy.
Dataset CIFAR10
# Label 40 250 4000
RegMixMatch (Mixup) 4.40$\pm$0.03 4.35$\pm$0.02 3.62$\pm$0.06
RegMixMatch 4.24$\pm$0.02 4.21$\pm$0.02 3.38$\pm$0.05
Table 7: Performance on other mixing strategy.
Pre-trained Backbones
Dataset CIFAR100 STL10 EuroSAT
# Label 200 400 40 100 20 40
VAT (Miyato et al. 2018) 31.49$\pm$1.33 21.34$\pm$0.50 18.45$\pm$1.47 10.69$\pm$0.51 26.16$\pm$0.96 10.09$\pm$0.94
Mean Teacher (Tarvainen and Valpola 2017) 35.47$\pm$0.40 26.03$\pm$0.30 18.67$\pm$1.69 24.19$\pm$10.15 26.83$\pm$1.46 15.85$\pm$1.66
MixMatch (Berthelot et al. 2019b) 38.22$\pm$0.71 26.72$\pm$0.72 58.77$\pm$1.98 36.74$\pm$1.24 24.85$\pm$4.85 17.28$\pm$2.67
ReMixMatch (Berthelot et al. 2019a) 22.21$\pm$2.21 16.86$\pm$0.57 13.08$\pm$3.34 7.21$\pm$0.39 5.05$\pm$1.05 5.07$\pm$0.56
UDA (Xie et al. 2020) 28.80$\pm$0.61 19.00$\pm$0.79 15.58$\pm$3.16 7.65$\pm$1.11 9.83$\pm$2.15 6.22$\pm$1.36
FixMatch (Sohn et al. 2020) 29.60$\pm$0.90 19.56$\pm$0.52 16.15$\pm$1.89 8.11$\pm$0.68 13.44$\pm$3.53 5.91$\pm$2.02
FlexMatch (Zhang et al. 2021) 26.76$\pm$1.12 18.24$\pm$0.36 14.40$\pm$3.11 8.17$\pm$0.78 5.17$\pm$0.57 5.58$\pm$0.81
CoMatch (Li, Xiong, and Hoi 2021) 35.08$\pm$0.69 25.23$\pm$0.50 15.12$\pm$1.88 9.56$\pm$1.35 5.75$\pm$0.43 4.81$\pm$1.05
SimMatch (Zheng et al. 2022) 23.78$\pm$1.08 17.06$\pm$0.78 11.77$\pm$3.20 7.55$\pm$1.86 7.66$\pm$0.60 5.27$\pm$0.89
SoftMatch (Chen et al. 2023) 22.67$\pm$1.32 16.84$\pm$0.66 13.55$\pm$3.16 7.84$\pm$1.72 5.75$\pm$0.62 5.90$\pm$1.42
FreeMatch (Wang et al. 2023) 21.40$\pm$0.30 15.65$\pm$0.26 12.73$\pm$3.22 8.52$\pm$0.53 6.50$\pm$0.78 5.78$\pm$0.51
RegMixMatch 19.26$\pm$0.56 15.55$\pm$0.31 10.11$\pm$3.20 7.10$\pm$0.66 4.25$\pm$0.77 3.61$\pm$0.75
Table 8: Error rates on CIFAR100, STL10 and EuroSAT datasets in USB. The best results are highlighted with Bold and the second-best results are highlighted with underline.
We also validate the performance of RegMixMatch on the USB (Wang et al. 2022) benchmark. USB adopts the pre-trained Vision Transformers (ViT) (Dosovitskiy 2020) instead of training ResNets (He et al. 2016) from scratch for CV tasks, thus achieving better performance with less training time.
We evaluate the performance of RegMixMatch on three datasets, including CIFAR-100 (Krizhevsky, Hinton et al. 2009), STL-10 (Coates, Ng, and Lee 2011), and Euro-SAT (Helber et al. 2019), under different labeled data conditions. The experimental results are compared against 11 established SSL algorithms, including VAT (Miyato et al. 2018), Mean Teacher (Tarvainen and Valpola 2017), MixMatch (Berthelot et al. 2019b), ReMixMatch (Berthelot et al. 2019a), UDA (Xie et al. 2020), FixMatch (Sohn et al. 2020), FlexMatch (Zhang et al. 2021), CoMatch (Li, Xiong, and Hoi 2021), SimMatch (Zheng et al. 2022), SoftMatch (Chen et al. 2023), and FreeMatch (Wang et al. 2023). For a fair comparison, we use the same setup presented in USB.
Table 8 shows the results of SSL algorithms on pre-trained backbones, where our RegMixMatch achieves comprehensive superiority. Notably, RegMixMatch still achieves significant improvement on the satellite dataset EuroSAT, where the accuracy of SSL algorithms is close to saturation.
ImageNet
Dataset ImageNet
# Label 1w 10w
Top-n acc Top-1 Top-5 Top-1 Top-5
FixMatch 53.61 75.96 71.53 90.36
FlexMatch 54.21 76.80 72.17 90.59
FreeMatch 54.69 77.02 72.57 90.97
RegMixMatch 58.35 81.10 73.66 91.89
Table 9: ImageNet accuracy results.
Following USB, we provide an evaluation on ImageNet of MAE pre-trained ViT-B (He et al. 2022), comparing RegMixMatch with FixMatch, Flexmatch, and FreeMatch. We train these algorithms using 10 labels per class and 100 labels per-class, i.e., a total of 10,000 labels and 100,000 labels respectively. We set the batch size for both labeled and unlabeled data to 64, use AdamW with a smaller learning rate of 3e-4 and a weight decay of 0.05. Other algorithmic hyper-parameters stay the same as USB. Our method, RegMixMatch, consistently achieves leading performance across various scenarios.